zarkam21:

The sum of two numbers is 48. If one third of one number is 5 greater than one sixth of another number, which of the following is the smaller number?

Shadow:

The way I think you would represent this is that let x and y be the two numbers. \[x + y = 48\] \[\frac{ 1 }{ 3 }x = \frac{ 1 }{ 6 }y + 5\]

Shadow:

Then it's just a system of equations and you solve

zarkam21:

how would you solve it if there are no like terms?

Shadow:

Isolate one variable in on equation to where it's like x = blah blah blah, then input it into the other equation then solve for y.

Shadow:

Then you input what you get for y back into the first equation, solving for x. Then you just take whichever of the numbers is smaller.

Shadow:

To answer, " which of the following is the smaller number?"

zarkam21:

I got 26

Shadow:

k now solve for the other number. Input it back into the first equation

zarkam21:

22 right?

Shadow:

Yes

zarkam21:

ok thank you so much

Shadow:

You're welcome